OpenStudy (anonymous):

what is the equation of the oblique asymptote h(x)=x^2-3x-4/x+1

OpenStudy (anonymous):

i got y=x-4

OpenStudy (anonymous):

@geerky42

OpenStudy (aum):

You have the correct answer.

OpenStudy (aum):

\[ h(x) = \frac{x^2-3x-4}{x+1} = \frac{(x-4)(x+1)}{x+1} = x-4 \text{ for }x \ne -1.\\ y = x- 4 \text{ is the equation of the oblique asymptote.} \]